OpenStudy (anonymous):

What is the standard error of the proportion if 80% of the 1,200 respondents in a yes/no survey answered "no"?

OpenStudy (anonymous):

A.3.46% B.2.31% C.1.15% D.2.58% E.1.29%

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@amistre64

OpenStudy (whpalmer4):

Okay, we did one just like this yesterday... http://openstudy.com/study#/updates/51dae0cee4b0d9a104d950f1

OpenStudy (whpalmer4):

standard error for a sample proportion, p = 0.8, n = 1200 \[SE = \sqrt{\frac{p(1-p)}{n}} = \sqrt{\frac{0.8(1-0.8)}{1200} }=\]